OpenStudy (anonymous):

find the minimum y-value on the graph of y=f(x). f(x)=x^2+9x

OpenStudy (anonymous):

take the derivative and set equal to 0 f'(x) = 2x + 9 = 0 2x = -9/2 x = -9/4 plug in to get y val y = (-9/4)^2 + 9(-9/4) y = 81/16 -81/4 = -243/16 = -15.1875

OpenStudy (anonymous):

or use the equation to find vertex -b/2a so -9/2(1) = -9/2 <- plug this in again to find y val

OpenStudy (anonymous):

so is the anwer -9 or 0 ?

OpenStudy (anonymous):

the answer is -15.1875

OpenStudy (anonymous):

\[x ^{2} +9x = (x+\frac{9}{2})^{2}-(\frac{9}{2})^{2}\] so the min is -81/4 = -20.25